# Fluid mechanics, Applied mathematics, Mathematical modelling

### v2

@@ 2.3 Truncation of Vertical Modes
We turn to modelling. Our first step is to introduce a Galerkin truncation in the vertical direction. Given the sinusoidal body force ${\bf f}$, corresponding sinusoidal laminar flow, and the stress-free boundary conditions, we use a sine-cosine basis and project onto the first two vertical modes (that is, the $y$-independent mode and $\sin(\beta y)$ or $\cos(\beta y)$ mode). The model large-scale flow and TKE are represented by
@@
The retained modes in (10) represent the minimal set necessary to describe a three-dimensional vector field, and DNS of Wf confirms that these modes are the most energetic in turbulent bands. It is therefore reasonable to base our model on this minimal set. While $v_{1}$ is small in magnitude, it is essential to the flow dynamics. For the TKE field, $q_{0}$ represents the $y$-averaged turbulence and $q_{1}$ accounts for an asymmetry between the upper and lower halves of the domain. This is important for capturing the so-called overhangs at the edges of turbulent bands, where turbulence is not vertically uniform due to the mean shear [48, 7, 25]. Substitution of (10) into (5b), followed by a Galerkin projection, yields evolution equations for the fields. However, before proceeding, we must first consider the turbulence closures.
@@ 2.4 Turbulence Closures
The turbulence closures are based on fully-resolved simulations of Wf, described in §2.1, as well as four-mode simulations of Wf from previous work [34]. Each of the four fluctuation correlation terms (6) - (9) has an expansion in vertical modes; however, it is only the pseudo-dissipation $\varepsilon$ that will explicitly require two vertical modes.
@@ 2.4.1 Reynolds Stress
We start with the Reynolds stress tensor. Only two components play a significant role in the dynamics, namely $-\langle u^{\prime}u^{\prime}\rangle$ and $-\langle u^{\prime}v^{\prime}\rangle$. Fully resolved simulations reveal only one dominant vertical mode in each of these tensor components – the zeroth vertical mode (invariant in $y$) for $-\langle u^{\prime}u^{\prime}\rangle$ and the first vertical mode for $-\langle u^{\prime}v^{\prime}\rangle$. The latter fact can be seen clearly in figure 4 of Chantry et al. (2016) [34]. It is assumed that the amplitudes of these terms depend only on the TKE and have the form:
@@
$\nabla\cdot{\bf\mathcal{R}}=\left(-\partial_{x}B(q)-A(q)\beta\sin(\beta y)\right){\bf e}_{x}+\left(\cos(\beta y)\partial_{x}A(q)\right){\bf e}_{y}.$ (13)
@@
We now neglect the $-\langle u^{\prime}u^{\prime}\rangle$ stress, and take $B(q)=0$ for the remainder of this work. The reason we have included it in the discussion to this point is because this stress accounts for the largest effect that we have measured in DNS but that we do not include in the present model. This stress couples the $u_{0}$ field to $q_{0}$ and $q_{1}$. Its associated Reynolds-stress force acts to reduce the amplitude of $u_{0}$, while its associated production term increases $q_{0}$ (its effect is small compared to other production terms). We have briefly investigated the role of this stress in the model and have found that including it does not alter the model dynamics in a significant way. Therefore, for simplicity, we neglect it in the present model. Future studies may wish to include it. We find that taking $B(q)$ linear in $q$ is a reasonable approximation.
@@
where $a$ and $\eta$ are parameters with $\eta$ small. Figure 3a compares $A(q_{0})$ with data from DNS. For $q_{0}\gg\eta,A(q_{0})\approx aq_{0}$, meaning that $\mathcal{R}_{12}$ is simply proportional to the TKE. For small $q_{0}$, that is, $q_{0}\ll\eta$, we have $A(q_{0})\approx aq_{0}^{2}/2\eta$, and $A(q_{0})$ goes to zero quadratically in $q_{0}$. This serves to cut off turbulent production at small $q_{0}$. The importance of this will be clear in the analysis of the model in §3.1. While $A(q_{0})$ underestimates the data for small $q_{0}$, and the Reynolds stresses do depend slightly on $Re$, we find that fitting the stresses to a more complicated functional form does not significantly alter the model dynamics. For simplicity, we choose not to account for these factors in the present model. Figure 4a compares DNS-measured values of the actual Reynolds-stress force across a turbulent band with closure (14), as well as with the corresponding force from model simulations of a turbulent band (figure 7a, $t=25350$).
Figure 3: Functional dependence of (a) the first mode amplitude of $-\langle u^{\prime}v^{\prime}\rangle$ on $q_{0}$, and (b) the vertically averaged pseudo-dissipation rate $\varepsilon_{0}$ on $q_{0}$. Data is from a DNS of a steady band in Wf at Re = 140, with averaging along the band and over 1000 time units. The black curve tracks the values of $-\langle u^{\prime}v^{\prime}\rangle$ or $\varepsilon_{0}$ with $q_{0}$ as the across-band direction is traversed. The maximum of $q_{0}$ occurs at the band centre, while small $q_{0}$ corresponds to the quasi-laminar regions between turbulent bands. Two lines appear because the averages extracted from DNS are not exactly symmetric about the band centre. See for example figure 4b for the variation of $\varepsilon_{0}$ across the band direction. In (a) the dashed line shows $A(q_{0})$ for the model and the vertical grey line highlights $q_{0}=\eta$, where $\eta=5\times 10^{-3}$. In (b) the dashed line (displaced vertically for clarity) represent best fits of the DNS data.
Figure 4: Comparison of Reynolds-stress force and pseudo-dissipation from DNS, model closures, and a model solution. Panel (a) shows the Reynolds-stress force in the $u_{1}$ mode. DNS results are obtained from a steady band in Wf at $Re=140$, with averaging along the band and over 1000 time units. The dashed curve shows the corresponding force using closure (14) with $q_{0}$ obtained from the DNS. The model curve (green) shows the force for a steady band in the model at $Re=75$. (See §2.6 for a discussion on the shift in $Re$.) Panels (b) and (c) show similar plots, but for the pseudo-dissipation rates $\varepsilon_{0}$ and $\varepsilon_{1}$, respectively.
@@ 2.4.3 Dissipation and Transport
As with $A(q)$, the even term in the pseudo-dissipation rate $\varepsilon_{0}$ depends only on $q_{0}$. DNS results show that the dependence of $\varepsilon_{0}$ on $q_{0}$ is approximately a power-law, with an exponent slightly greater than one (figure 3b). In keeping with our goal of model simplicity, we model this dependence as linear. Tests confirm that a linear approximation is adequate (figure 4b). DNS results also support that the coefficient of proportionality depends on $Re$ approximately as $Re^{-1}$ throughout the transition region. Hence, our closure for $\varepsilon_{0}$ is
@@
where $c$ is a model parameter. A similarly simple closure exists for $\varepsilon_{1}(q)$, that is $\varepsilon_{1}(q_{1})=\kappa q_{1}$, where $\kappa$ does not depend on $Re$ (figure 4c).
Moving on to the transport term, DNS results and model tests suggest that only the zeroth vertical mode of the transport term plays a significant role in the turbulence budget. Apart from this simplification in vertical mode number, we invoke the gradient-diffusion hypothesis [47, §10.3] which states that,
@@
We have again used the fact that, since ${\bf T}$ does not depend on $y$, it will be only a function of $q_{0}$. The turbulent diffusivity is defined in a standard way as $\nu_{T}\propto q^{2}_{0}/\varepsilon_{0}$. (See Pope [47, §10.3].) Since we have already assumed that $\varepsilon_{0}=cq_{0}/Re$, we can then substitute this into the turbulent diffusivity arriving at,
$\nu_{T}(q_{0};Re)=d\,Re\,q_{0}.$ (17)
@@ 2.5 Quasi-static approximation for the $q_{1}$ dynamics
[…]
@@ 2.6 Parameter fits
Model parameters have been calibrated starting from a DNS of a turbulent band in Wf at $Re=140$. From measured Reynolds-stress amplitudes, turbulent dissipation, and turbulent diffusion, we obtained a first estimate for model parameters. We subsequently refined values to ensure model simulations match reasonably well those of the DNS, as shown in figure 5, albeit with a necessary shift in $Re$ (discussed momentarily). We summarize here the most relevant aspects of the parameter fitting.
@@
The Reynold stress and dissipation from the DNS of Wf are seen in figure 3. The model parameters $a$ and $c$ associated with these effects, (14) and (15), are particularly important as they dictate the onset of sustained turbulence in the model. (See §3.1.) We attempted to tweaked $a$ and $c$ so that sustained turbulence first appears at Reynolds numbers close to those observed in Wf. However, matching these Reynolds numbers results in the formation of sharp gradients in the model large-scale flow, which also leads to instability and unphysical phenomena. We therefore settled on values of $a$ and $c$ such that model turbulent bands occur around $Re\simeq 75$. These $Re$ values are roughly half of the corresponding values in Wf [34]. This shift to lower $Re$ is a common feature in modal truncations [34, 33]. Indeed, turbulent bands occur in the range $50\lesssim Re\lesssim 90$ in a four-vertical-mode truncation of Wf [34, 12], so it is sensible to aim for such Reynolds numbers in our model.
@@
After calibrating $a$ and $c$, we performed simulations of turbulent bands using the model equivalent of the tilted domain technique [23], much like the DNS used for calibration. Assuming bands are tilted at $24^{\circ}$ to the streamwise direction and are invariant in the along-band direction, the model reduces to a system of one-dimensional (1D) PDEs in the across-band coordinate. We performed a series of simulations and observed how the turbulent bands’ width, amplitude, and disappearance to uniform flow depended on $Re$ and other model parameters. From these runs, we modified the rest of the parameters, in particular the turbulent diffusivity coefficient $d$, so as to best match the results found in the DNS.
@@
Throughout our experimentation in the tilted 1D domain, we observed steady bands and their transition scenario (figure 8 discussed below) to be extremely robust to parameter changes. In full 2D model simulations, we also observed initial band formation to be a very robust phenomenon. However, experimenting with parameter values in 2D simulations also revealed various instabilities in the along-band direction and cases where the large-scale-flow amplitude is unrealistically large. Neglecting the Reynolds stress $\langle u^{\prime}u^{\prime}\rangle$, which plays a significant role in the budget of $u_{0}$, could be an issue, but tests including $B(q_{0})$ did not solve the issue of band instabilities. We believe that ultimately the issue is that the truncation of vertical modes results in too little dissipation in the large-scale flow, especially in the zero modes $u_{0}$ and $w_{0}$. Even after choosing parameter values that lower Reynolds numbers in the model, we found it necessary to increase the value of the drag coefficient $\alpha$ for the zero modes. This solves the issues of instabilities and the amplitude of the large-scale flow. We let $\alpha_{0}$ denote the drag coefficient for the zero modes and in practice use $\alpha_{0}=3\alpha$. The introduction of $\alpha_{0}$ is our final modelling assumption.
@@ 2.7 Summary
In summary, the model is represented by six fields, $u_{0},u_{1},v_{1},w_{0},w_{1},q_{0}$, corresponding to the mode amplitudes of large-scale flow and TKE. The evolution equations of these fields in $(x,z,t)$ are obtained by substituting expansion (10) into the Reynolds average equations (5b), using model closures for ${\bf\mathcal{R}}$, $\mathcal{P}$, $\varepsilon$, and T, using a quasi-static approximation for $q_{1}$, imposing incompressibility, and applying a Galerkin projection. The six evolution equations, together with the model closures and parameter value are stated in full in appendix A.

### v3

@@ 2.3 Truncation of Vertical Modes
We turn to modelling. Our first step is to introduce a Galerkin truncation in the vertical direction. Severe truncations, in both the wall-normal and streamwise directions, have been shown to successfully reproduce turbulent statistics at both low[manneville2011modelling, chantry2016turbulent, chantry_universal, seshasayanan2015laminar, lagha2007largescaleflow, lagha2007modeling] and high[Thomas2014, Constantinou2014, Farrell2017, Cavalieri2022, Minnick2024] Reynolds numbers in a variety of settings such as plane Couette and plane Poiseuille flows. Given the sinusoidal body force $\bf f$}, corresponding sinusoidal laminar flow, and the stress-free boundary conditions, we use a sine-cosine basis and project onto the first two vertical modes (that is, the $y$-independent mode and $\sin(\beta y)$ or $\cos(\beta y)$ mode). The model large-scale flow and TKE are represented by
@@
The retained modes in (10) represent the minimal set necessary to describe a three-dimensional vector field, and DNS of Wf confirms that these modes are the most energetic in turbulent bands. It is therefore reasonable to base our model on this minimal set. While $v_{1}$ is small in magnitude, it is essential to the flow dynamics. For the TKE field, $q_{0}$ represents the $y$-averaged turbulence and $q_{1}$ accounts for an asymmetry between the upper and lower halves of the domain. This is important for capturing the so-called overhangs at the edges of turbulent bands, where turbulence is not vertically uniform due to the mean shear [coles1966progress, barkley2007mean, duguet2013oblique]. Substitution of (10) into (5b), followed by a Galerkin projection, yields evolution equations for the fields. However, before proceeding, we must first consider the turbulence closures.
@@ 2.4 Turbulence Closures
The turbulence closures are based on DNS of Wf, described in §2.1, as well as four-mode simulations of Wf from previous work [chantry2016turbulent]. Each of the four fluctuation correlation terms (6) - (9) has an expansion in vertical modes; however, it is only the pseudo-dissipation $\varepsilon$ that will explicitly require two vertical modes.
@@ 2.4.1 Reynolds Stress
We start with the Reynolds stress tensor. The DNS reveals that only two components play a significant role in the dynamics, namely $\langle u^{\prime}v^{\prime}\rangle$ and $\langle u^{\prime}u^{\prime}\rangle$. In our truncated space of vertical modes, $\langle u^{\prime}v^{\prime}\rangle$ has non-zero projection onto only the first vertical mode $\cos(\beta y)$. The truncation is supported by DNS results for a steady band in Wf at $Re=140$, from which the root-mean-square (RMS) magnitude of the first vertical mode of $\langle u^{\prime}v^{\prime}\rangle$ is, respectively, 10.2, 7.5, and 67.7 times larger than the second, third, and fourth vertical modes. The $\langle u^{\prime}u^{\prime}\rangle$ component projects onto both the zeroth (constant in $y$) and first vertical modes, but from DNS, the zeroth mode is dominant (5.2 times larger than the first vertical mode for a steady band in Wf at $Re=140$) and we consider only the zeroth mode. It is assumed that the amplitudes of these Reynolds stress components depend only on the TKE and have the form:
@@
$\nabla\cdot{\bf\mathcal{R}}=\left(-A(q)\beta\sin(\beta y)-\partial_{x}B(q)\right){\bf e}_{x}+\left(\cos(\beta y)\partial_{x}A(q)\right){\bf e}_{y}.$ (13)
@@
We now neglect the $\langle u^{\prime}u^{\prime}\rangle$ stress, and take $B(q)=0$ for the remainder of this work. The reason we have included it in the discussion to this point is because this stress accounts for the largest effect that we have measured in DNS but that we do not include in the present model. This stress couples the $u_{0}$ field to $q_{0}$ and $q_{1}$. Its associated Reynolds-stress force acts to reduce the amplitude of $u_{0}$, while its associated production term increases $q_{0}$ (its effect is small compared to other production terms). We have briefly investigated the role of this stress in the model and have found that including it does not alter the model dynamics in a significant way. Therefore, for simplicity, we neglect it in the present model. Future studies may wish to include it. We find that taking $B(q)$ linear in $q$ is a reasonable approximation.
@@
where $a$ and $\eta$ are parameters with $\eta$ small. Figure 3a compares $A(q_{0})$ with data from DNS. For $q_{0}\gg\eta,A(q_{0})\approx aq_{0}$, meaning that $\mathcal{R}_{12}$ is simply proportional to the TKE. For small $q_{0}$, that is, $q_{0}\ll\eta$, we have $A(q_{0})\approx aq_{0}^{2}/2\eta$, and $A(q_{0})$ goes to zero quadratically in $q_{0}$. This serves to cut off turbulent production at small $q_{0}$. The importance of this will be clear in the analysis of the model in §3.1. While $A(q_{0})$ underestimates the data for small $q_{0}$, and the Reynolds stresses do depend slightly on ReRe, we find that fitting the stresses to a more complicated functional form does not significantly alter the model dynamics. For simplicity, we choose not to account for these factors in the present model. Figure 4a compares DNS-measured values of the actual Reynolds-stress force across a turbulent band with closure (14), as well as with the corresponding force from model simulations of a turbulent band (figure 7a, $t=25350$).
Figure 3: Functional dependence of (a) the first mode amplitude of $\langle u^{\prime}v^{\prime}\rangle$ on $q_{0}$, and (b) the vertically averaged pseudo-dissipation rate $\varepsilon_{0}$ on $q_{0}$. Data is from a DNS of a steady band in Wf at $Re=140$, with averaging along the band and over 1000 time units. The black curve tracks the values of $\langle u^{\prime}v^{\prime}\rangle$ or $\varepsilon_{0}$ with $q_{0}$ as the across-band direction is traversed. The maximum of $q_{0}$ occurs at the band centre, while small $q_{0}$ corresponds to the quasi-laminar regions between turbulent bands. Two lines appear because the averages extracted from DNS are not exactly symmetric about the band centre. See for example figure 4b for the variation of $\varepsilon_{0}$ across the band direction. In (a) the dashed line shows $A(q_{0})$ for the model and the vertical grey line highlights $q_{0}=\eta$, where $\eta=5\times 10^{-3}$. In (b) the dashed line (displaced vertically for clarity) represent best fits of the DNS data.
Figure 4: Comparison of Reynolds-stress force and pseudo-dissipation from DNS, model closures, and a model solution. Panel (a) shows the Reynolds-stress force in the $u_{1}$ mode. DNS results are obtained from a steady band in Wf at $Re=140$, with averaging along the band and over 1000 time units. The dashed curve shows the corresponding force using closure (14) with $q_{0}$ obtained from the DNS. The model curve (green) shows the force for a steady band in the model at $Re=75$. (See §2.6 for a discussion on the shift in ReRe.) Panels (b) and (c) show similar plots, but for the pseudo-dissipation rates $\varepsilon_{0}$ and $\varepsilon_{1}$, respectively.
@@ 2.4.3 Dissipation and Transport
As with $A(q)$, the even term in the pseudo-dissipation rate $\varepsilon_{0}$ depends only on $q_{0}$. DNS results show that the dependence of $\varepsilon_{0}$ on $q_{0}$ is approximately a power-law, with an exponent slightly greater than one (figure 3b). In keeping with our goal of model simplicity, we model this dependence as linear. Tests confirm that a linear approximation is adequate (figure 4b). DNS results also support that the coefficient of proportionality depends on ReRe approximately as $Re^{-1}$ throughout the transition region. Hence, our closure for $\varepsilon_{0}$ is
@@
where $c$ is a model parameter. A similarly simple closure exists for $\varepsilon_{1}(q)$, that is $\varepsilon_{1}(q_{1})=\kappa q_{1}$, where $\kappa$ does not depend on ReRe (figure 4c).
Turning to the transport term $\bf T$}, DNS results and model tests indicate that only the zeroth vertical mode of the transport term plays a significant role in the turbulence budget, and thus in our model we ignore all but the zeroth vertical mode for this term. We invoke the gradient-diffusion hypothesis [pope2000turbulent, §10.3] which states that,
@@
This is the simple physical assertion that (due to microscopic velocity and pressure fluctuations) there is a flux of $q_{0}$ down the gradient of $q_{0}$, with a transport coefficient, turbulent diffusivity $\nu_{T}$, itself depending on the $q_{0}$. The turbulent diffusivity is defined in a standard way to be proportional $q^{2}_{0}/\varepsilon_{0}$. (See Pope [pope2000turbulent, §10.3].) Given the form for $\varepsilon_{0}=cq_{0}/Re$, equation (15), the turbulent diffusivity thus has the form
$\nu_{T}(q_{0};Re)=d\,Re\,q_{0},$ (17)
where $d$ is a model parameter.
@@ 2.5 Quasi-static approximation for the $q_{1}$ dynamics
$\cancel{\frac{\partial q_{1}}{\partial t}}+u_{1}\partial_{x}q_{0}+w_{1}\partial_{z}q_{0}+\cancel{u_{0}\partial_{x}q_{1}}+\cancel{w_{0}\partial_{z}q_{1}}=-(2\alpha+\beta^{2}/Re)q_{1}-\kappa q_{1},$ (18)
@@ 2.6 Parameter fits
Model parameters have been calibrated starting from a DNS of a turbulent band in Wf at $Re=140$. From measured Reynolds-stress amplitudes, turbulent dissipation, and turbulent diffusion, we obtained a first estimate for model parameters. We subsequently refined values to ensure model simulations match reasonably well those of the DNS, as shown in figure 5, albeit with a necessary shift in ReRe (discussed momentarily). We summarize here the most relevant aspects of the parameter fitting.
@@
The Reynold stress and dissipation from the DNS of Wf are seen in figure 3. The model parameters $a$ and $c$ associated with these effects, (14) and (15), are particularly important as they dictate the onset of sustained turbulence in the model. (See §3.1.) We attempted to tweak $a$ and $c$ so that sustained turbulence first appears at Reynolds numbers close to those observed in Wf. However, matching these Reynolds numbers results in the formation of sharp gradients in the model large-scale flow, which also leads to instability and unphysical phenomena. We therefore settled on values of $a$ and $c$ such that model turbulent bands occur around $Re\simeq 75$. These ReRe values are roughly half of the corresponding values in Wf [chantry2016turbulent]. This shift to lower ReRe is a common feature in modal truncations [chantry2016turbulent, manneville2011modelling]. Indeed, turbulent bands occur in the range $50\lesssim Re\lesssim 90$ in a four-vertical-mode truncation of Wf [chantry2016turbulent, chantry_universal], so it is sensible to aim for such Reynolds numbers in our model. Our study will focus on Reynolds numbers in the range $60$ to $100$, where the model exhibits the spatiotemporal dynamics of transitional turbulence. Outside the range the model produces only either fully laminar or fully turbulent solutions.
@@
After calibrating $a$ and $c$, we performed simulations of turbulent bands using the model equivalent of the tilted domain technique [barkley2005computational], much like the DNS used for calibration. Assuming bands are tilted at $24^{\circ}$ to the streamwise direction and are invariant in the along-band direction, the model reduces to a system of one-dimensional (1D) PDEs in the across-band coordinate, which we term the ‘1D tilted domain’. We performed a series of simulations and observed how the turbulent bands’ width, amplitude, and disappearance to uniform flow depended on ReRe and other model parameters. From these runs, we modified the rest of the parameters, in particular the turbulent diffusivity coefficient $d$, so as to best match the results found in the DNS. We did not attempt a more sophisticated parameter optimization beyond what is described here.
@@
Throughout our experimentation in the tilted 1D domain, we observed steady bands and their transition scenario to be extremely robust to parameter changes. In the 2D model simulations, we also observed initial band formation to be a very robust phenomenon. However, experimenting with parameter values in 2D simulations also revealed various instabilities in the along-band direction and cases where the large-scale-flow amplitude is unrealistically large. Neglecting the Reynolds stress $\langle u^{\prime}u^{\prime}\rangle$, which plays a significant role in the budget of $u_{0}$, could be an issue, but tests including $B(q_{0})$ did not solve the issue of band instabilities. We believe that ultimately the issue is that the truncation of vertical modes results in too little dissipation in the large-scale flow, especially in the zero modes $u_{0}$ and $w_{0}$. Even after choosing parameter values that lower Reynolds numbers in the model, we found it necessary to increase the value of the drag coefficient $\alpha$ for the zero modes. This solves the issues of instabilities and the amplitude of the large-scale flow. We let $\alpha_{0}$ denote the drag coefficient for the zero modes and in practice use $\alpha_{0}=3\alpha$. The introduction of $\alpha_{0}$ is our final modelling assumption.
@@ 2.7 Summary
[…]
$\frac{\partial u_{0}}{\partial t}$ $+(u_{0}\partial_{x}+w_{0}\partial_{z})u_{0}+\frac{1}{2}(u_{1}\partial_{x}+\beta v_{1}+w_{1}\partial_{z})u_{1}=-\partial_{x}p_{0}+\left(\frac{1}{Re}\nabla^{2}_{H}-\alpha_{0}\right)u_{0},$ (20a)
$\frac{\partial u_{1}}{\partial t}$ $+(u_{0}\partial_{x}+w_{0}\partial_{z})u_{1}+(u_{1}\partial_{x}+w_{1}\partial_{z})u_{0}=-\partial_{x}p_{1}+\left(\frac{1}{Re}(\nabla^{2}_{H}-\beta^{2})-\alpha\right)u_{1}$
$\hskip 170.71652pt+f-\uwave{\beta A(q_{0})},$ (20b)
$\frac{\partial v_{1}}{\partial t}$ $+(u_{0}\partial_{x}+w_{0}\partial_{z})v_{1}=-\beta p_{1}+\frac{1}{Re}(\nabla^{2}_{H}-\beta^{2})v_{1}+\uwave{\partial_{x}A(q_{0})},$ (20c)
$\frac{\partial w_{0}}{\partial t}$ $+(u_{0}\partial_{x}+w_{0}\partial_{z})w_{0}+\frac{1}{2}(u_{1}\partial_{x}+\beta v_{1}+w_{1}\partial_{z})w_{1}=-\partial_{z}p_{0}+\left(\frac{1}{Re}\nabla^{2}_{H}-\alpha_{0}\right)w_{0},$ (20d)
$\frac{\partial w_{1}}{\partial t}$ $+(u_{0}\partial_{x}+w_{0}\partial_{z})w_{1}+(u_{1}\partial_{x}+w_{1}\partial_{z})w_{0}=-\partial_{z}p_{1}+\left(\frac{1}{Re}(\nabla^{2}_{H}-\beta^{2})-\alpha\right)w_{1},$ (20e)
$=\frac{1}{2}\uwave{A(q_{0})}[\beta u_{1}+\partial_{x}v_{1}]-2\alpha q_{0}-\uwave{\varepsilon_{0}(q_{0};Re)}+\frac{1}{Re}\nabla^{2}_{H}q_{0}+\uwave{\nabla_{H}\cdot\left(\nu_{T}(q_{0};Re)\nabla_{H}q_{0}\right)},$ (20f)
where $f=\alpha+\beta^{2}/Re$, $\beta=\pi/2$, and $\nabla_{H}=(\partial_{x},\partial_{z})$. The terms coming from the turbulence closure are highlighted by wavy underline. The only forcing of the large-scale flow by Reynolds stresses appears only via two terms. All other terms in the momenta equations arises just from modelling by Ekman friction and a Galerkin projection of the Navier-Stokes equations onto the retained modes.
The equations are closed with the following,
Equations (21) constitute the turbulence closures, expressing the dependence of Reynolds stress, pseudo-dissipation, and turbulent transport on turbulence amplitude $q_{0}$. Equation (22) follows from the quasi-static approximation for the $q_{1}$ field.
Model parameters have the values
$a=0.3,\,\eta=5\times 10^{-3},\,c=9.65,\,d=0.05,\,\kappa=0.09,\,\alpha=0.01,\,\alpha_{0}=0.03.$ (23)
The parameters $a$ and $\eta$ control the Reynolds stress closure, with $a$ setting its amplitude and $\eta$ setting the turbulence level below which the turbulence production rapidly decreases. Parameter $c$ sets the amplitude of turbulent dissipation associated with $q_{0}$; $d$ is the coefficient for the diffusivity of $q_{0}$. Parameter $\kappa$ sets the turbulent dissipation for $q_{1}$, which for the model affects the amplitude of $q_{1}$ via the quasi-static approximation. Finally, $\alpha$ is the Ekman drag coefficient and $\alpha_{0}$ is the modified Ekman drag coefficient for the zeroth vertical modes.